OpenStudy (anonymous):

I will give medals what is the value of z so that -9 and 9 are both the same solutions of x^2 + z = 103 A. -22 B. 3 C. 22 D. 184

OpenStudy (ikram002p):

if 9 ,-9 is for x then ovs f(9)=f(-9) :D

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

is it 22

OpenStudy (ikram002p):

well make f(9)=0 i thought somone was solving it O.O

OpenStudy (anonymous):

im confused

OpenStudy (ikram002p):

since 9,-9 arw solutions then f(9)=0

OpenStudy (anonymous):

but that's not a option

OpenStudy (ikram002p):

yep its option f(x)=x^2+z-103 f(9)=9^2+z-103=0 z=-22

OpenStudy (ikram002p):

yep